OpenStudy (anonymous):

B is between A and C. Find BC if AB = 3, BC = 4x + 1, AC = 8.

OpenStudy (anonymous):

Here AC is the full line segment and AB is a part of it also BC = AC - AB = 8 - 3 = 5 => 4x + 1 = 5 => 4x = 5-1 => x = 4/4 = 1

OpenStudy (anonymous):

if u r happy with the answer, pls click GOOD ANSWER button for me :) Pls also become my fan as I need one more fan to complete the number of 50 fans ( :})